A Bibliography of Numerical Analysis Publications, 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.04

Title word cross-reference

Advances [APE22]. Algebra [BMR23]. Analysis [APE22].

Data [APE22, BHK20].

Emphasizing [APE22]. Engineers [PW20]. Essential [PW20].

Foundations [BHK20].

History [BMR23].

Interval [APE22].

Journey [BMR23].

Linear [BMR23].

Mathematics [PW20].

Numerical [APE22, BMR23].

Science [BHK20]. Scientists [PW20].

References

Allahviranloo:2022:ANA

[APE22] Tofigh Allahviranloo, Witold Pedrycz, and Armin Esfandiari. Advances in Numerical Analysis Emphasizing Interval Data. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2022. ISBN 1-00-054025-1 (e-book), 1-00-054031-6 (ePUB), 1-00-321817-2 (ebook), 1-03-211043-0 (hardcover). 224 (est.) pp. LCCN QA297 .A55 2022.

Blum:2020:FDS

[BHK20] Avrim Blum, John Hopcroft, and Ravi Kannan. Foundations of Data Science. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-48506-5 (hardcover), 1-108-75552-6 (e-book). viii + 424 pp. LCCN QA76 .B5675 2020.

Brezinski:2023:JTH

[BMR23] Claude Brezinski, Gérard A. Meurant, and Michela Redivo Zaglia. A Journey Through the History of Numerical Linear Algebra, volume 183 of Other titles in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2023. ISBN 1-61197-722-3 (hardcover), 1-61197-723-1 (ebook). xix + 792 pp. LCCN QA184.2.

Pence:2020:EMEd

[PW20] T. J. (Thomas J.) Pence and I. S.

(Indrek S.) Wichman. Essential Mathematics for Engineers and Scientists. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-67135-7. xix + 736 pp. LCCN QA37.3 .P46 2020.